OpenStudy (anonymous):

How to integrate (x+3)/sqrt(x)?

OpenStudy (accessdenied):

Have you tried breaking it up into two separate fractions (x/sqrt(x) + 3/sqrt(x)) and integrating each term individually?

OpenStudy (accessdenied):

I believe if we change sqrt(x) into its exponent representation x^(1/2), we have the opportunity of using power rule on each separate term's integral. x/sqrt(x) = x/(x^1/2) = x^(1-1/2) = x^(1/2)

OpenStudy (anonymous):

Okay. Wow. I was completely overthinking it. Thanks for your help!

OpenStudy (accessdenied):

You're welcome! :)